A man stands before a large wall at a distance of 50.0 m and claps his hands at regular intervals. Initially, the interval is large. He gradually reduces the interval and fixes it at a value when the echo of a clap merges with the next clap. If he has to clap 10 times during every 3 seconds, find the
velocity of sound in air.

Grade:12

1 Answers

Arun
6 years ago
 He has to clap 10 times in 3 seconds. So time interval between two clap = (3/10 second). So the time taken go the wall = (3/2 × 10) = 3/20 seconds. = 333 m/s.